OpenStudy (anonymous):

48.0g piece of metal is dropped into 50.0mL of water in a graduated cylinder. The water level rises to 62.4mL. What is the density of the metal?

OpenStudy (lgbasallote):

first find the volume \[V = V_F - V_1\] your final volume is 62.4 mL and initial volume is 50.0 mL so... \[\implies V = 62.4 - 50.0\] simplify... \[\implies V = 12.4 \; \text{mL}\] now to get density use the formula \[\rho = \frac MV\] here, mass is 48.0 g so density is \[\implies \rho = \frac{48.0 \; g}{12.4 ~ mL}\] use a calculator to solve dies that help?

OpenStudy (lgbasallote):

@jess.evans are you there?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

sorry i just joined and i have no idea how to use this

OpenStudy (lgbasallote):

it's really just simple. you ask something you're confused about...we try to teach you how to solve what you're confused about. you give medals to people who helped you the most. you close your question when you finished. then post a new question. cycle repeats. then, if you see a question by others that you know you can teach, feel free to teach

OpenStudy (anonymous):

i signed up for 6 online classes so im completely overwhelmed im a single mom taking double full time school and works full time so as you can imagine i need help lol

OpenStudy (lgbasallote):

use your calculator and solve what i wrote....